OpenStudy (anonymous):

"Rita's father is seven times as old as Rita is. One year ago, he was nine times as old as Rita was. Find Rita's present age." Anyone know how?

OpenStudy (amistre64):

maybe; if im quick i can get it done b4 library closes :)

OpenStudy (amistre64):

f = 7r f-1 = 9r 7r -1 = 9r -1 = 2r r = -1/2....doesnt seem right lol

OpenStudy (amistre64):

f = 7r f-1 = 9(r-1) might be more better

OpenStudy (amistre64):

7r -1 = 9r -9 8 = 2r r = 4 f = 28 ; r = 4

OpenStudy (anonymous):

Rita's father: x Rita: y i) 7y = x ii) 9*(y-1) = (x-1) i) 7y = x ii) 9y - x - 8 = 0 i -> ii) 9y - 7y - 8 = 0 y = 4 i) x = 7y x = 28

OpenStudy (amistre64):

why add confusion by naming variables that have no relation to the elements that you are examining? never made sense to me... Rita = r ; Rita father = f gives you a visual on what you are equating.... That said; that variables dont care what they are called :) so x and y is just as valid....